# HCVP: Leveraging Hierarchical Contrastive Visual Prompt for Domain Generalization

### v1

@@ V-A1 Datasets
We evaluate on five benchmark datasets in DG [13, 32]. PACS [59]: encapsulates four distinctive visual domains: Art Paintings, Cartoon, Photos, and Sketches; contains 9,991 instances across 7 classes, each being a (3, 224, 224)-dimensional image. VLCS [60]: an amalgamation of four photographic domains: Caltech-101, LabelMe, SUN09, and VOC2007; includes 10,729 instances in five classes, each a (3, 224, 224)-dimensional image. Office-Home [61]: comprises four domains: Art, Clipart, Product, and Real-World Images, with 15,588 images in 65 classes, each of dimension (3, 224, 224). TerraIncognita [62]: features wildlife images from four camera trap locations (L100, L38, L43, L46), each representing a distinct domain; has 24,788 (3, 224, 224)-dimensional images across 10 classes. CelebA [63]: focuses on “hair color” classification against “gender” as a spurious attribute [64, 32]. Our subset has 27,040 images in three environments, akin to the Colored MNIST dataset’s correlation shift [65, 32].
@@ V-A2 Baselines
We compare HCVP with two categories of DG methods: DG Algorithms include ERM [66], IRM [65], MLDG [67], MMD [43], DANN [7], CDANN [11], VREx [68], MTL [69], SagNet [70], RSC [71], IB_ERM, IB_IRM [72], SWAD [36], MIRO [37] and POEM [35]; Tuning Protocols include full fine-tuning, linear probing (LP)—a frozen backbone with a tuned linear head [19], partial-k (fine-tuning last k layers). We also acknowledge related visual prompt methods like DoPrompt [27]. A direct comparison with CSVPT [28] is not feasible, as its implementation is not open-sourced.
@@ V-A3 Implementation Details
[…]
TABLE I: Performance comparison of HCVP with existing DG algorithms and tuning protocols on five DG datasets. The best and second-best results are marked in Red and Blue respectively.
Methods PACS VLCS OfficeHome TerraIncognita CelebA Avg
DG Algorithms
ERM 89.3 80.0 82.5 54.7 84.6 78.2
IRM 82.9 78.1 73.6 38.9 82.1 71.1
MLDG 89.9 79.8 82.5 52.9 85.0 78.0
MMD 88.4 79.9 82.1 53.0 85.7 77.8
DANN 84.2 79.0 80.1 47.8 85.5 75.3
CDANN 84.3 78.4 80.4 49.4 84.5 75.4
VREx 87.3 79.4 80.3 53.1 84.5 76.9
MTL 87.2 80.1 80.7 55.0 84.5 77.5
SagNet 83.5 80.1 77.6 48.3 83.8 74.7
RSC 89.8 79.9 82.5 55.5 85.0 78.5
IB_ERM 82.2 78.5 71.6 33.6 83.4 69.9
IB_IRM 81.4 78.3 58.4 37.7 82.6 67.7
SWAD 90.1 79.3 79.7 53.0 84.0 77.2
MIRO 83.6 77.4 66.0 47.7 81.4 71.2
POEM 84.3 76.1 73.3 39.3 83.7 71.3
Tuning Protocols
Full 89.3 80.0 82.5 54.7 84.6 78.2
LP 71.6 76.9 75.6 35.2 71.0 66.1
Partial_1 74.5 79.4 76.1 36.7 79.7 69.3
Partial_2 76.7 79.9 76.8 38.4 81.1 70.6
Partial_4 80.5 80.3 77.7 38.8 84.6 72.4
DoPrompt 89.8 80.3 82.7 54.5 85.3 78.5
Our Method
HCVP 90.2 81.1 82.5 55.1 86.6 79.1
@@ V-B Main Results
In our evaluation, HCVP is benchmarked against established DG algorithms and various tuning protocols. The comparative results are presented in Table I. For a more detailed analysis, particularly focusing on performance across each unseen domain, refer to Tables II, III, IV, and V. Our comparison spans two primary types of distribution shifts: diversity shift observed in the PACS, VLCS, OfficeHome and TerraIncognita datasets, and the spurious shift in the CelebA dataset [32].
V-B1 DG Algorithms
@@ V-B2 Tuning Protocols
We outline tuning protocols for adapting visual foundation models to distribution shifts in downstream tasks. As shown in Table I, full fine-tuning, akin to ERM, yields a robust average accuracy of 78.2%. In contrast, less computationally intensive methods such as linear probing and gradual unfreezing (Partial_1, Partial_2, Partial_4) achieve relatively lower average accuracies, almost all falling below 72%. This trend clearly illustrates a correlation between the number of tuned parameters and the average performance under distribution shifts. In this context, our HCVP integrates domain and task-specific information into the model, by augmenting visual prompts to the ViT backbone. This enhancement not only aligns with the observed trend of increased performance with broader parameter tuning but also contributes to the model’s robustness. For example, compared to the best tuning protocol—full fine-tuning—our HCVP outperforms it on all five datasets. This advantage becomes particularly pronounced when facing substantial distribution shifts. For example, when adapting to the sketch domain on the PACS dataset, denoted by "normal-$\rightarrow$ Sketch ", the performance is 81.30% compared to 75.13%, as shown in Table II. Similarly, when encountering spurious distribution shifts on the CelebA dataset, the performance is 86.6% versus 84.6%.
@@ V-B3 Significance of Incremental Gains in DG
In the field of DG, even marginal performance improvements carry considerable significance. This is largely attributed to the inherent complexity and the unpredictable nature of generalizing effectively to unseen domains. The previous large-scale evaluations [13, 32, 14] indicate that no single DG algorithm consistently outperforms over ERM across all benchmark datasets. However, our HCVP method demonstrates this remarkable capability. Therefore, the incremental gains, such as the 1% improvement reported by HCVP, should be recognized for their substantial value.
@@ V-C Quantifying Domain-Invariance
We initially explored the possibility of directly measuring the conditional mutual information $I(Z,D|Y)$ to assess our model’s capacity to learn invariant features [54]. This measurement is consistent with Eq. 2 to evaluate the hypothesis that the incorporation of $P$ leads to the reduction of $I(Z,D|Y)$. However, given the high-dimensional nature of the latent representation $Z$, we encountered significant computational challenges in reliably estimating this metric. To circumvent this issue, we opted to measure inter-domain feature distances as an alternative metric. This approach is grounded in the theory that lower distances across domains indicate more domain-invariant feature learning [75, 7]. In essence, closer feature representations across domains suggest that the model is not overfitting to domain-specific characteristics and is, instead, capturing more generalizable aspects of the data. Specifically, we regard the last domain in each dataset as the unseen domain. We then train both ERM and HCVP models on this unseen domain and save the iterations of the models that demonstrate the best performance. Using the saved best model, we calculate the average inter-domain distance by utilizing the latent representations generated by the models.
@@ V-D Ablation Study
In this section, we perfrom ablation study to evaluate the individual contribution of each loss in our HCVP model. As shown in Table VI, where “P, V, O, T, C” represents PACS, VLCS, OfficeHome, TerraIncognita and CelebA respectively, the complete model "HCVP (full)" achieves the highest accuracy on each dataset and the best agerage accuracy of 79.10. Removing each loss sequentially leads to "HCVP w/o $\mathcal{L}_{PCL}$" and "HCVP w/o $\mathcal{L}_{CCI}$", which exhibit slightly reduced accuracies of 78.59 and 78.21, respectively. This underscores the importance of these losses to the model’s performance. Removing both losses results in "HCVP w/o $\mathcal{L}_{CCI}$ and $\mathcal{L}_{PCL}$", further reducing the accuracy to 77.69.
TABLE VI: This ablation study evaluates the individual contributions of various losses in the HCVP model across four datasets. "HCVP (full)" refers to the complete model. We sequentially remove each loss to create "HCVP w/o $\mathcal{L}_{PCL}$" and "HCVP w/o $\mathcal{L}_{CCI}$". Removing both losses results in "HCVP w/o $\mathcal{L}_{CCI}$ and $\mathcal{L}_{PCL}$", the vanilla HCVP variant with only the classification loss $\mathcal{L}_{cls}$. Values are shown with two decimal place; columns P, V, O, T, C correspond to the PACS, VLCS, OfficeHome, TerraIncognita, and CelebA datasets.
@@ V-E1 The t-SNE Visualizations of Visual Features
Figure 4 illustrates the t-SNE [76] visualizations of visual features in our HCVP, for the last unseen domain on the PACS, VLCS, OfficeHome, CelebA and TerraIncognita datasets. Overall, the visual embeddings of HCVP are clearly clustered in accordance with their class labels across five datasets. This clear clustering is indicative of the robustness and effectiveness of our HCVP model in capturing invariant features for task prediction. In generalizing to the last unseen domain on PACS, namely “Sketch”, where the test scenario may diverge largely from our pretraining and training features in “Art, Photo, Cartoon”, an overlap between several classes, such as dogs with elephants and horses, is observed in Figure 4 (a). Despite this, our HCVP still shows the best generalization ability, as evidenced in Tables I and II.
@@ V-E2 The t-SNE Visualizations of Visual Prompts
To investigate the relationship between visual prompts with domain and task specific characteristics, we also leverage t-SNE visualizations, as visual prompts typically lack clear meanings [25, 26]. As illustrated in Figure 5, our visualization study on the PACS dataset reveals: (a) domain prompt vectors corresponding to the same domains are clustered together, these vectors are distinctly separate from other domains, validating that they carry domain-level characteristics; (b) task prompts are similarly clustered by specific task labels. These findings collectively affirm that the visual prompts generated by our method effectively capture domain-level and task-specific characteristics, contributing to HCVP’s superior performance in DG tasks.
@@ V-F Hyperparameter Tuning
[…]

### v2

@@ V Experiments
TABLE I: Performance comparison of HCVP with existing DG algorithms and tuning protocols on five DG datasets. The best and second-best results are marked in Red and Blue respectively. OH and TI represent OfficeHome and TerraIncognita datasets. Algorithms are ordered according to their publication years for ease of reference.
Methods PACS VLCS OfficeHome TerraIncognita CelebA Average
DG Algorithms
ERM (1998) [vapnik1998statistical] 89.3 80.0 82.5 54.7 84.6 78.2
DANN (2016) [ganin2016domain] 84.2 79.0 80.1 47.8 85.5 75.3
MLDG (2018) [li2018learning] 89.9 79.8 82.5 52.9 85.0 78.0
MMD (2018) [li2018domain] 88.4 79.9 82.1 53.0 85.7 77.8
CDANN (2018) [li2018deep] 84.3 78.4 80.4 49.4 84.5 75.4
IRM (2019) [arjovsky2019invariant] 82.9 78.1 73.6 38.9 82.1 71.1
SagNet (2019) [nam2019reducing] 83.5 80.1 77.6 48.3 83.8 74.7
RSC (2020) [huang2020self] 89.8 79.9 82.5 55.5 85.0 78.5
VREx (2021) [krueger2021out] 87.3 79.4 80.3 53.1 84.5 76.9
MTL (2021) [blanchard2021domain] 87.2 80.1 80.7 55.0 84.5 77.5
IB_ERM (2021) [ahuja2021invariance] 82.2 78.5 71.6 33.6 83.4 69.9
IB_IRM (2021) [ahuja2021invariance] 81.4 78.3 58.4 37.7 82.6 67.7
SWAD (2021) [cha2021swad] 90.1 79.3 79.7 53.0 84.0 77.2
Transfer (2021) [zhang2021quantifying] 83.0 78.8 75.8 48.8 80.2 73.3
MIRO (2022) [cha2022domain] 83.6 77.4 66.0 47.7 81.4 71.2
EQRM (2022) [eastwood2022probable] 90.0 78.7 82.8 54.0 83.7 77.8
CausIRL-MMD (2022) [chevalley2022invariant] 89.0 80.8 82.5 53.1 84.0 77.9
POEM (2023) [jo2023poem] 84.3 76.1 73.3 39.3 83.7 71.3
SAGM (2023) [wang2023sharpness] 90.4 80.3 82.5 54.8 82.7 78.2
Tuning Protocols
Full (1998) [vapnik1998statistical] 89.3 80.0 82.5 54.7 84.6 78.2
Linear Probing (2022) [kumar2022fine] 71.6 76.9 75.6 35.2 71.0 66.1
Partial_1 (2022)  [lee2022surgical] 74.5 79.4 76.1 36.7 79.7 69.3
Partial_2 (2022)  [lee2022surgical] 76.7 79.9 76.8 38.4 81.1 70.6
Partial_4 (2022)  [lee2022surgical] 80.5 80.3 77.7 38.8 84.6 72.4
DoPrompt (2022) [zheng2022prompt] 89.8 80.3 82.7 54.5 85.3 78.5
Our Method
HCVP 90.2 81.1 82.5 55.1 86.6 79.1
@@ V-A1 Datasets
We evaluate on five benchmark datasets in DG [gulrajani2020search, ye2022ood]. PACS [li2017deeper]: encapsulates four distinctive visual domains: Art Paintings, Cartoon, Photos, and Sketches; contains 9,991 instances across 7 classes, each being a (3, 224, 224)-dimensional image. VLCS [fang2013unbiased]: an amalgamation of four photographic domains: Caltech-101, LabelMe, SUN09, and VOC2007; includes 10,729 instances in five classes, each a (3, 224, 224)-dimensional image. Office-Home [venkateswara2017deep]: comprises four domains: Art, Clipart, Product, and Real-World Images, with 15,588 images in 65 classes, each of dimension (3, 224, 224). TerraIncognita [beery2018recognition]: features wildlife images from four camera trap locations (L100, L38, L43, L46), each representing a distinct domain; has 24,788 (3, 224, 224)-dimensional images across 10 classes. CelebA [liu2015deep]: focuses on “hair color” classification against “gender” as a spurious attribute [sagawa2019distributionally, ye2022ood]. Our subset has 27,040 images in three environments, akin to the Colored MNIST dataset’s correlation shift [arjovsky2019invariant, ye2022ood].
@@ V-A2 Baselines
We compare HCVP with two categories of DG methods: DG Algorithms include ERM [vapnik1998statistical], IRM [arjovsky2019invariant], MLDG [li2018learning], MMD [li2018domain], DANN [ganin2016domain], CDANN [li2018deep], VREx [krueger2021out], MTL [blanchard2021domain], SagNet [nam2019reducing], RSC [huang2020self], IB_ERM, IB_IRM [ahuja2021invariance], SWAD [cha2021swad], MIRO [cha2022domain] and POEM [jo2023poem]; Tuning Protocols include full fine-tuning, linear probing (LP)—a frozen backbone with a tuned linear head [kumar2022fine], partial-k (fine-tuning last k layers). We also acknowledge related visual prompt methods like DoPrompt [zheng2022prompt]. A direct comparison with CSVPT [li2022learning] is not feasible, as its implementation is not open-sourced.
@@ V-A3 Implementation Details
We adopt unified settings for all methods: ViT-B/16 backbone [^2] See ‘vit_base_patch16_224’ in https://github.com/huggingface/pytorch-image-models/blob/main/timm/models/vision_transformer.py pre-trained on ImageNet-1k and finetuned on ImageNet-21k [dosovitskiy2021an, deng2009imagenet]; AdamW optimizer [loshchilov2017decoupled] with learning rate 1ee-5 and weight decay 0.01; 80%:20% train-validation split, batch size of 32, 3000 steps per trial; averages reported over 3 runs with different seeds. Training-domain validation strategy is used for model selection. We use PyTorch to implement all experiments on NVIDIA A100-40GB and A5000-24GB GPUs. Specifically, the PMN comprises a series of linear layers mirroring the ViT’s architecture, with each layer autonomously modulating domain and task prompts for precise alignment with its corresponding ViT layer.
@@ V-B Main Results
In our evaluation, HCVP is benchmarked against established DG algorithms and various tuning protocols. The comparative results are presented in Table I. For a more detailed analysis, particularly focusing on performance across each unseen domain, refer to Tables II, III, IV, and V. Our comparison spans two primary types of distribution shifts: diversity shift observed in the PACS, VLCS, OfficeHome and TerraIncognita datasets, and the spurious shift in the CelebA dataset [ye2022ood].
@@ V-B2 Tuning Protocols
We outline tuning protocols for adapting visual foundation models to distribution shifts in downstream tasks. As shown in Table I, full fine-tuning, akin to ERM, yields a robust average accuracy of 78.2%. In contrast, less computationally intensive methods such as linear probing and gradual unfreezing (Partial_1, Partial_2, Partial_4) achieve relatively lower average accuracies, almost all falling below 72%. This trend clearly illustrates a correlation between the number of tuned parameters and the average performance under distribution shifts. In this context, our HCVP integrates domain and task-specific information into the model, by augmenting visual prompts to the ViT backbone. This enhancement not only aligns with the observed trend of increased performance with broader parameter tuning but also contributes to the model’s robustness. For example, compared to the best tuning protocol—full fine-tuning—our HCVP outperforms it on all five datasets. This advantage becomes particularly pronounced when facing substantial distribution shifts. For example, when adapting to the sketch domain on the PACS dataset, denoted by $\rightarrow$ Sketch ", the performance is 81.30% compared to 75.13%, as shown in Table II. Similarly, when encountering spurious distribution shifts on the CelebA dataset, the performance is 86.6% versus 84.6%.
@@ V-B3 Significance of Incremental Gains in DG
In the field of DG, even marginal performance improvements carry considerable significance. This is largely attributed to the inherent complexity and the unpredictable nature of generalizing effectively to unseen domains. The previous large-scale evaluations [gulrajani2020search, ye2022ood, wiles2022a] indicate that no single DG algorithm consistently outperforms over ERM across all benchmark datasets. However, our HCVP method demonstrates this remarkable capability. Therefore, the incremental gains, such as the 1% improvement reported by HCVP, should be recognized for their substantial value.
@@ V-C Quantifying Domain-Invariance
We initially explored the possibility of directly measuring the conditional mutual information $I(Z,D|Y)$ to assess our model’s capacity to learn invariant features [li2022invariant]. This measurement is consistent with Eq. 2 to evaluate the hypothesis that the incorporation of $P$ leads to the reduction of $I(Z,D|Y)$. However, given the high-dimensional nature of the latent representation $Z$, we encountered significant computational challenges in reliably estimating this metric. To circumvent this issue, we opted to measure inter-domain feature distances as an alternative metric. This approach is grounded in the theory that lower distances across domains indicate more domain-invariant feature learning [ben2010theory, ganin2016domain]. In essence, closer feature representations across domains suggest that the model is not overfitting to domain-specific characteristics and is, instead, capturing more generalizable aspects of the data. Specifically, we regard the last domain in each dataset as the unseen domain. We then train both ERM and HCVP models on this unseen domain and save the iterations of the models that demonstrate the best performance. Using the saved best model, we calculate the average inter-domain distance by utilizing the latent representations generated by the models.
@@ V-D Ablation Study
In this section, we perfrom ablation study to evaluate the individual contribution of each loss in our HCVP model. As shown in Table VI, where “P, V, O, T, C” represents PACS, VLCS, OfficeHome, TerraIncognita and CelebA respectively, the complete model "HCVP (full)" achieves the highest accuracy on each dataset and the best agerage accuracy of 79.10. Removing each loss sequentially leads to "HCVP w/o $\mathcal{L}_{PCL}$" and "HCVP w/o $\mathcal{L}_{CCI}$", which exhibit slightly reduced accuracies of 78.59 and 78.21, respectively. This underscores the importance of these losses to the model’s performance. Removing both losses results in "HCVP w/o $\mathcal{L}_{CCI}$ and $\mathcal{L}_{PCL}$", further reducing the accuracy to 77.69. Intriguingly, our expanded analysis revealed that removing either $\mathcal{L}_{domain}$ or $\mathcal{L}_{task}$ detrimentally impacts the model’s performance more significantly than omitting both in the form of $\mathcal{L}_{PCL}$. This finding suggests a complex interplay between $\mathcal{L}_{domain}$ and $\mathcal{L}_{task}$, where their combined presence bolsters the model’s domain generalization capability more than their individual contributions might imply.
TABLE VI: This ablation study evaluates the individual contributions of various losses in the HCVP model across four datasets. "HCVP (full)" refers to the complete model. We sequentially remove each loss to create "HCVP w/o $\mathcal{L}_{task}$", "HCVP w/o $\mathcal{L}_{domain}$" "HCVP w/o $\mathcal{L}_{PCL}$" and "HCVP w/o $\mathcal{L}_{CCI}$". Removing $\mathcal{L}_{PCL}$ and $\mathcal{L}_{CCI}$ results in "HCVP w/o $\mathcal{L}_{CCI}$ and $\mathcal{L}_{PCL}$", the vanilla HCVP variant with only the classification loss $\mathcal{L}_{cls}$. Values are shown with two decimal place; columns P, V, O, T, C correspond to the PACS, VLCS, OfficeHome, TerraIncognita, and CelebA datasets.
@@
HCVP w/o $\mathcal{L}_{domain}$ 88.88 80.38 81.91 54.62 85.42 78.24
HCVP w/o $\mathcal{L}_{task}$ 88.63 80.92 81.63 55.25 85.50 78.38
@@ V-E1 The t-SNE Visualizations of Visual Features
Figure 4 illustrates the t-SNE [Maaten2008VisualizingDU] visualizations of visual features in our HCVP, for the last unseen domain on the PACS, VLCS, OfficeHome, CelebA and TerraIncognita datasets. Overall, the visual embeddings of HCVP are clearly clustered in accordance with their class labels across five datasets. This clear clustering is indicative of the robustness and effectiveness of our HCVP model in capturing invariant features for task prediction. In generalizing to the last unseen domain on PACS, namely “Sketch”, where the test scenario may diverge largely from our pretraining and training features in “Art, Photo, Cartoon”, an overlap between several classes, such as dogs with elephants and horses, is observed in Figure 4 (a). Despite this, our HCVP still shows the best generalization ability, as evidenced in Tables I and II.
@@ V-E2 The t-SNE Visualizations of Visual Prompts
To investigate the relationship between visual prompts with domain and task specific characteristics, we also leverage t-SNE visualizations, as visual prompts typically lack clear meanings [zhou2022learning, jia2022visual]. As illustrated in Figure 5, our visualization study on the PACS dataset reveals: (a) domain prompt vectors corresponding to the same domains are clustered together, these vectors are distinctly separate from other domains, validating that they carry domain-level characteristics; (b) task prompts are similarly clustered by specific task labels. These findings collectively affirm that the visual prompts generated by our method effectively capture domain-level and task-specific characteristics, contributing to HCVP’s superior performance in DG tasks.
@@ V-F Hyperparameter Tuning
We investigate the effects of two loss weights, $\lambda_{PCL}$ and $\lambda_{CCI}$, on the VLCS and OfficeHome datasets. We explore a wide range of values for both weights, specifically $\{$0.001, 0.01, 0.1, 0.5, 1.0}\} for $\lambda_{PCL}$ and $\{$0.01, 0.1, 0.3, 0.6, 1.0}\} for $\lambda_{CCI}$. Given that the original value of $\mathcal{L}_{PCL}$ is almost four times larger than $\mathcal{L}_{\text{cls}}$, a large weight for $\lambda_{PCL}$ could impact classification. As depicted in Figure 6, when $\lambda_{PCL}$ is set to 0.1, our HCVP achieves optimal performance. Additionally, we find that 1.0 serves as an effective weight for $\mathcal{L}_{CCI}$.
@@
V-G Efficiency Analysis
The efficiency of DG algorithms is pivotal, balancing computational demands with performance [gulrajani2020search]. Table VII presents an efficiency analysis comparing our HCVP model against both traditional and recent state-of-the-art DG methods, evaluating total trainable parameters, gradient parameters within the featurizer, GPU memory usage, step time, and average performance across five datasets.
HCVP demonstrates an increment in total trainable parameters to 90,131,719, indicative of the additional hierarchical prompts and networks engineered to bolster domain generalization. Despite this increase, HCVP maintains competitive efficiency with a GPU memory usage of 13.463 GB—comparable to baseline methods with 13.177 GB—and a step time of 1.249 seconds, showcasing notable efficiency, particularly against the 2.404s of DoPrompt and 11.014s of Transfer model.
Crucially, HCVP secures the highest average accuracy of 79.1, outstripping close contenders such as RSC and DoPrompt. This underscores the effectiveness of our model’s complexity, balanced against a modest rise in resource consumption, to achieve superior domain generalization. In essence, this analysis confirms the thoughtful trade-off between model complexity and computational efficiency in our HCVP.
TABLE VII: Efficiency analysis of SOTA and recent DG algorithms. “Tot. params” indicates the total number of trainable parameters, “Grad. in feat.” denotes the gradient parameters within the featurizer, “GPU (GB)” represents the GPU memory usage in gigabytes, “Step time (s)” specifies the average time per training step in seconds, and “Avg. Perf.” reflects the average accuracy under five datasets.
Methods Tot. params Grad. in feat. GPU (GB) Step time (s) Avg. Perf.
ERM 85,804,039 85,798,656 13.177 1.102 78.2
MMD 85,804,039 85,798,656 12.947 1.117 77.8
RSC 85,804,039 85,798,656 13.179 1.170 78.5
DoPrompt 88,212,499 85,798,656 13.847 2.404 78.5
SAGM 85,802,501 85,798,656 13.177 2.001 78.2
EQRM 85,804,039 85,798,656 11.857 1.795 77.8
Transfer 85,809,422 85,798,656 11.959 11.014 73.3
CausIRL-MMD 85,804,039 85,798,656 11.857 0.992 77.9
HCVP 90,131,719 90,126,336 13.463 1.249 79.1